• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name V. Sulkosky

    Log entry time 09:54:26 on November 5,2009

    Entry number 297575

    keyword=S2m Gain Matching

    L-HRS S2m gain matching mostly completed. It appears that the left
    side channels 1 and 2 are swapped. Ramesh and Kai will investigate.
    It was probably swapped at the patch panel on the second floor.
    All channels are aligned to about 35 channels.

    Figures 1 and 2 show channels 0-5 and 6-11 respectively.

    FIGURE 1

    FIGURE 2